The function, p(d)=1+d/33, gives the pressure, in atmospheres (atm) at a depth d in the sea ( d is in feet). Note that p(0)=1 atm, p(33)=2, and so on. find the pressure at 50 feet.

Answer by edjones(8007)

p(50)= 1 + 50/33
= 33/33 + 50/33
=83/33 atm
.
Ed